OpenStudy (anonymous):

1/4 of the people own cars that are blue and 1/3 of the people have brown eyes. What is the probability that a person selected at random has a blue car and brown eyes

OpenStudy (anonymous):

C: event that people have cars that are blue. B: event that people have brown eyes. Because C and B are independence, \[P(B cupC)=P(B).P(C)=\] (1/3)(1/4)=1/12